OpenStudy (deadshot):

Simplify quantity x plus 3 over 9 all over quantity x plus 12 over 6

OpenStudy (deadshot):

\[\frac{ \frac{ x+3 }{ 9 } }{ \frac{ x+12 }{ 6 } }\]

OpenStudy (campbell_st):

you are dividing by a fraction, so flip the denominator and multiply \[\frac{x +3}{9} \times \frac{6}{x + 12} = \] hope this helps...

OpenStudy (deadshot):

so, it would be \[\frac{ 6x+18 }{ 9x+108 }\] right?

OpenStudy (campbell_st):

yes... but you can also remove a common factor of 3... divide each term by 3 gives \[\frac{2x +9}{3x + 36}\]

OpenStudy (deadshot):

Okay, Thanks!